# Adversarial Flow Models

### v1

@@ Abstract
We present adversarial flow models, a class of generative models that unifies adversarial models and flow models. Our method supports native one-step or multi-step generation and is trained using the adversarial objective. Unlike traditional GANs, where the generator learns an arbitrary transport plan between the noise and the data distributions, our generator learns a deterministic noise-to-data mapping, which is the same optimal transport as in flow-matching models. This significantly stabilizes adversarial training. Also, unlike consistency-based methods, our model directly learns one-step or few-step generation without needing to learn the intermediate timesteps of the probability flow for propagation. This saves model capacity, reduces training iterations, and avoids error accumulation. Under the same 1NFE setting on ImageNet-256px, our B/2 model approaches the performance of consistency-based XL/2 models, while our XL/2 model creates a new best FID of 2.38. We additionally show the possibility of end-to-end training of 56-layer and 112-layer models through depth repetition without any intermediate supervision, and achieve FIDs of 2.08 and 1.94 using a single forward pass, surpassing their 2NFE and 4NFE counterparts.
@@ 1 Introduction
Flow matching [49, 80] is a class of generative models that achieves the current state of the art in multiple domains, including image generation [11, 77, 3] and video generation [86, 38, 63, 76, 13], etc. It formulates the generation problem as the transport of samples from a prior distribution to the data distribution. A probability flow is established by interpolating the data and the prior samples, and a neural network learns the gradient of the flow. At inference time, the sample is iteratively transported by querying the network for gradients, incurring a high computational cost.
Recent methods accelerate the generation process by formulating the network to directly predict farther positions on the flow instead of the instantaneous gradients. This can be achieved either by distilling from a pre-trained flow-matching model [71, 50], or by training from scratch with consistency objectives as a new class of generative models [81]. However, despite the goal of single-step or few-step generation, consistency-based models must still be trained on all timesteps of the flow for consistency propagation. This wastes model capacity and incurs error accumulation. A more critical problem is that networks in fewer steps have less capacity to predict the exact transformations of the more-step targets, so using pointwise-matching or even moment-matching losses causes some extent of blurriness. For these reasons, many state-of-the-art few-step generation models still rely on distributional matching methods, especially adversarial training, for final refinement [47, 6].
@@
Figure 1: Models trained on 1D Gaussian mixture. (a) GAN learns an arbitrary transport plan. (b) Flow matching learns a deterministic transport plan but have large discretization errors on low sampling steps. (c) Adversarial flow models support any-step training and generation with a deterministic optimal transport plan.
@@
Adversarial training originates from generative adversarial networks (GANs) [15]. It itself is a standalone class of generative models that supports single-step generation. However, adversarial training from scratch often faces stability issues. A handful of recent works exploring adversarial training from scratch adopt non-standard architectural designs [21, 26, 108, 23]. Some further rely on the assistance of additional frozen feature networks [26, 23]. When we switch to a standard transformer architecture [85], the training simply diverges.
@@
As shown in Fig. 1, we discover that one of the key reasons GANs are hard to train is that the adversarial objective alone does not present a single optimization target. This is markedly different from other established objectives, such as flow-matching, which has a unique ground-truth probability flow predetermined by the interpolation function, and autoregressive modeling, which has ground-truth token probabilities predetermined by the training corpus. In GANs, the generator is tasked to transport samples from the prior to the data distribution, but the adversarial objective only enforces the matching of the data distributions without constraining the transport plan. Therefore, there exist infinite valid transport plans that the generator may pick, influenced by the weight initialization and the stochastic training process. This causes optimization difficulties as the generator keeps drifting during training.
@@
In this paper, we propose adversarial flow models, a class of generative models that unifies the adversarial and flow families of models under the same framework. Our model is trained by the adversarial objective. Therefore, it naturally supports single-step training and generation without wasting capacities on learning the intermediate timesteps needed by consistency methods. Our model also belongs to the flow family of models. Therefore, it has the same deterministic transport mapping for training stability, and it naturally supports multi-step training and generation. Our method can be trained on standard transformer architectures without modifications, which opens the door for wider adoption.
@@
On ImageNet-256px, our B/2 model approaches the performance of consistency-based XL/2 models due to the conservation of modeling capacity, while our XL/2 model creates a new best FID of 2.38 under the same 1NFE setting with guidance. Our method also enables fully end-to-end training of 56-layer and 112-layer 1NFE models through depth repetition, and achieves FIDs of 2.08 and 1.94, surpassing the 2NFE and 4NFE counterparts. Furthermore, in the same no-guidance setting, we show that the adversarial objective can even outperform flow matching due to better distribution matching properties.
@@ 2 Related Works
Early distillation works train few-step student models to match the prediction of the teacher flow model [71, 50, 72, 94]. Consistency model (CM) [81, 79] proposes the use of self-consistency constraint and supports standalone training as a new class of generative models. Various follow-up works improve upon CM. Notably, sCM [52] extends the consistency constraint to continuous time to minimize discretization error. iMM [107] incorporates moment matching. Shortcut [12] redefines the boundary condition to allow jumping between arbitrary timesteps. MeanFlow [14] again extends Shortcut to continuous time. However, these methods typically still generate slightly blurry results on large-scale text-to-image/video tasks [54], so distributional matching methods, such as adversarial training [46, 66, 44, 47, 48, 53, 75, 74, 87, 37, 27, 92, 6] and score distillation [97, 96, 75, 53, 55, 105], are often incorporated in practice.
@@
Early research in GANs [15] has developed many techniques to achieve success on domain-specific datasets [65, 100, 28, 29, 31, 32, 30]. BigGAN [4] and StyleGAN-XL [73] have further scaled it to ImageNet [70]. However, GANs have fallen out of favor because of their training instability and limited scalability. A handful of works exploring large-scale text-to-image generation using GANs still employ a convolutional architecture with complex designs [26, 108]. GANs with transformer architectures [85] have been challenging to scale up [24, 40, 22]. Until recently, R3GAN [21] simplifies the adversarial formulation and pushes the state of the art in the ImageNet-64 benchmark using a convolutional architecture. GAT [23] further extends it to a latent transformer architecture. These works revitalize interest in adversarial training. However, GAT still employs a non-standard transformer architecture and uses the assistance of a pre-trained feature network. Our work provides a framework that unifies adversarial models with flow models and improves upon the training stability of adversarial methods.
@@ 3.1 Adversarial Training Preliminaries
Our method is built on top of generative adversarial networks (GANs) [15], where a generator $G:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ aims to transport samples $z$ from a prior distribution $\mathcal{Z}$, $\emph{e.g}.\hbox{}$ Gaussian, to samples $x$ from the data distribution $\mathcal{X}$, and a discriminator $D:\mathbb{R}^{n}\rightarrow\mathbb{R}^{1}$ aims to differentiate real samples from the generated ones. The adversarial optimization involves a minimax game where $D$ is trained to maximize differentiation while $G$ is trained to minimize the differentiation by $D$. We adopt the relativistic objective [25] as it has better properties for the loss landscape [82] and achieves the current state of the art [21, 23]:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x}\left[f(D(x)-D(G(z)))\right],$ (1)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x}\left[f(D(G(z))-D(x))\right],$ (2)
where $f(\cdot)=-\log(\mathrm{sigmoid}(\cdot))$.
Additionally, gradient penalties $R_{1}$ and $R_{2}$ [69] are added on $D$. They prevent $G$ from being pushed away from equilibrium [57, 21] and impose a constraint on the Lipschitz constant of $D$ [16]. Directly computing the gradient penalties requires expansive double backward and twice differentiation, so we use finite difference approximation [47]:
$\mathcal{L}^{D}_{\mathrm{r1}}$ $=\mathbb{E}_{x}\left[\|\triangledown_{x}D(x)\|^{2}_{2}\right]$ (3)
@@
$\mathcal{L}^{D}_{\mathrm{r2}}$ $=\mathbb{E}_{z}\left[\|\triangledown_{G(z)}D(G(z))\|^{2}_{2}\right]$ (5)
@@
where $\epsilon$ is set to a small value 0.01. We only compute the penalties on 25% of the samples of a batch and observe no performance degradation.
To prevent the discriminator logits from drifting unboundedly in the relativistic setting, a logit-centering penalty is added, similar to prior work [28]:
$\small\mathcal{L}^{D}_{\mathrm{cp}}=\mathbb{E}_{z,x}\left[\|D(x)+D(G(z))\|^{2}_{2}\right].$ (7)
Therefore, the final GAN loss is:
@@
where $\lambda_{\mathrm{gp}}$ is a searched hyperparameter controlling the gradient penalty scale of both $R_{1}$ and $R_{2}$, and $\lambda_{\mathrm{cp}}$ is universally set to 0.01.
The expectation is obtained through Monte Carlo approximation over minibatches of data during training. The $G$ and $D$ are updated in alternation. For conditional generation, the condition $c$ is given to both networks as $G(z,c)$ and $D(x,c)$. These are abbreviated in writing.
@@ 3.2 Single-step Adversarial Flow Models
The GAN objective above only enforces $G(z)$ to match the data distribution. However, there are infinite transport plans from $z$ to $x$, and the model is free to learn any arbitrary one. Our proposed adversarial flow models learn a deterministic transport plan, which is the same transport plan as flow-matching models with the linear interpolation function. This allows for the existence of a single solution, prevents generator drifting, and stabilizes training.
In flow matching, the ground-truth probability flow exists prior to model training and is deterministically defined by the data distribution, the prior distribution, the interpolation function, and the loss objective. It is known that when linear interpolation and the squared distance loss are used, this combination establishes a transport plan that minimizes the squared Wasserstein-2 ($\mathbf{W_{2}^{2}}$) distance between the prior and data distributions [49].
Formally, let $\bm{\pi}(x,z)$ be a valid transport plan, and $\mathbf{c}(x,z)$ be a cost function, flow matching with linear interpolation and squared distance objective results in the transport plan $\bm{\pi}(x,z)$ that minimizes the total transport cost where the cost function $\mathbf{c}(x,z)=\|x-z\|^{2}_{2}$ by the definition of $\mathrm{W}_{2}^{2}$:
$\small\min_{\pi}\int_{\mathcal{X}\times\mathcal{Z}}\mathbf{c}(x,z)\,d\bm{\pi}(x,z).$ (10)
Therefore, for adversarial flow models, we first restrict the prior distribution to have the same dimensionality $n$ as the data distribution, $\emph{i.e}.\hbox{},x,z\in\mathbb{R}^{n}$ and $G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This is required by all flow-based models and does not reduce the generality of our method. Then we substitute $\pi(x,z)$ with our generator $G(z)$:
$\small\min_{G}\int_{\mathcal{Z}}\mathbf{c}(G(z),z)\,dG(z).$ (11)
Our goal is to find a $G$ that is simultaneously a valid transport plan and a plan that minimizes the total transport cost under our defined cost function. Since the adversarial objective already enforces the validity of the transport plan, we only need to add an additional optimal transport (OT) loss on $G$. The loss minimizes the expectation of $\mathbf{c}(G(z),z)$ for all $z$:
$\small\boxed{\mathcal{L}^{G}_{\mathrm{ot}}=\mathbb{E}_{z}\left[\frac{1}{n}\|G(z)-z\|^{2}_{2}\right].}$ (12)
The objectives in adversarial flow models (AF) become:
@@
Figure 2: The effect of using different $\lambda_{\textrm{ot}}$ scale. $\lambda_{\textrm{ot}}=0$ is equivalent to GANs. $\lambda_{\textrm{ot}}$ being too small fails to escape local minima. $\lambda_{\textrm{ot}}$ being too large forces identity output.
In theory, even an infinitesimal $\lambda_{\mathrm{ot}}$ will break the symmetry in the loss landscape and create a unique global minimum. But in practice, if the value is too small, the optimization may get stuck in local minima, and if the value is too large, it forces the model to output identity, hurting distribution matching, as illustrated in Fig. 2. Therefore, we adopt a schedule to decrease $\lambda_{\mathrm{ot}}$ over time during training.
Unlike consistency-based methods, our model does not need to be trained on other timesteps of the probability flow and can be trained for native one-step generation. This saves model capacity, reduces training iterations, and avoids error propagation. Furthermore, for the one-step training case, this formulation completely removes the hyperparameter of choosing timestep sampling and weighting. Our one-step model also completely avoids teacher-forcing.

### v2

@@ Abstract
We present adversarial flow models, a class of generative models that belongs to both the adversarial and flow families. Our method supports native one-step and multi-step generation and is trained with an adversarial objective. Unlike traditional GANs, in which the generator learns an arbitrary transport map between the noise and data distributions, our generator is encouraged to learn a deterministic noise-to-data mapping. This significantly stabilizes adversarial training. Unlike consistency-based methods, our model directly learns one-step or few-step generation without having to learn the intermediate timesteps of the probability flow for propagation. This preserves model capacity and avoids error accumulation. Under the same 1NFE setting on ImageNet-256px, our B/2 model approaches the performance of consistency-based XL/2 models, while our XL/2 model achieves a new best FID of 2.38. We additionally demonstrate end-to-end training of 56-layer and 112-layer models without any intermediate supervision, achieving FIDs of 2.08 and 1.94 with a single forward pass and surpassing the corresponding 28-layer 2NFE and 4NFE counterparts with equal compute and parameters. The code is available at this repository.
Generative models, Adversarial training, Flow models
@@ 1 Introduction
Flow matching (Lipman et al., 2022) is a generative method that has achieved state-of-the-art performance across multiple domains. It frames generation as transporting samples from a prior distribution to the data distribution. A probability flow is established by interpolating between data samples and prior samples, and a neural network learns the gradient field of this flow. At inference time, each sample is transported iteratively by querying the network for gradients, incurring a high computational cost.
Recent methods accelerate generation by training networks to predict distant positions along the flow rather than instantaneous gradients. This can be achieved either by distilling from a pre-trained flow-matching model (Salimans and Ho, 2022; Liu et al., 2022) or by training from scratch with consistency objectives, forming a new class of generative models (Song et al., 2023). However, even when targeting single-step or few-step generation, consistency-based models must still be trained across all timesteps to propagate consistency. This consumes model capacity and introduces error accumulation. Furthermore, models operating in fewer steps have less capacity to predict the exact transformations of targets produced with more steps, so pointwise matching or even moment-matching losses can lead to some degree of blurriness. For these reasons, many state-of-the-art few-step generation models still rely on distributional matching methods, especially adversarial training, for final refinement (Lin et al., 2025a; Chen et al., 2025).
@@
Figure 1: Models trained on a 1D Gaussian mixture. (a) GANs learn arbitrary transport maps. (b) Flow matching learns a deterministic transport map but have large discretization errors on low sampling steps. (c) Adversarial flow models support any-step training and generation with a deterministic optimal transport map.
@@
Adversarial training originates from generative adversarial networks (GANs) (Goodfellow et al., 2014). It is itself a standalone class of generative models that supports single-step generation. However, adversarial training from scratch often suffers from stability issues. Recent works exploring adversarial training from scratch adopt non-standard architectures (Huang et al., 2024; Zhu et al., 2025). Some further rely on additional frozen feature networks (Kang et al., 2023; Hyun et al., 2025). When we switch to a standard transformer architecture (Vaswani et al., 2017), training simply diverges.
@@
As shown in Figure 1, we find that one of the key reasons GANs are difficult to train is that the adversarial objective alone does not define a single optimization target. This differs markedly from other established objectives, such as flow matching, which has a unique ground-truth probability flow determined by the interpolation function, and autoregressive modeling, which has ground-truth token probabilities determined by the training corpus. In GANs, the generator is tasked with transporting samples from the prior to the data distribution, but the adversarial objective only enforces matching between the data distributions without constraining the transport map. Therefore, there are infinitely many valid transport maps that the generator may choose, depending on the weight initialization and the stochastic training process. This creates optimization difficulties because the generator keeps drifting during training.
@@
In this paper, we propose adversarial flow models, a class of generative models that belongs to both the adversarial and flow families. Our models are trained with an adversarial objective, which naturally supports single-step training and generation without consuming model capacity to learn the intermediate timesteps required by consistency methods. At the same time, they belong to the flow family. Like flow matching, they learn a deterministic transport map, which improves training stability and naturally generalizes to multi-step training and generation. Our method can be trained on standard transformer architectures without modification, opening the door to wider adoption.
@@
On ImageNet-256px, our B/2 model approaches the performance of consistency-based XL/2 models because it preserves modeling capacity, while our XL/2 model achieves a new best FID of 2.38 under the same 1NFE setting. Our method also enables fully end-to-end training of 56-layer and 112-layer 1NFE models through depth repetition without any intermediate supervision, achieving FIDs of 2.08 and 1.94 and surpassing their 28-layer 2NFE and 4NFE counterparts.
@@ 2 Related Works
Early distillation works train few-step student models to match the predictions of a teacher flow model (Salimans and Ho, 2022; Liu et al., 2022; Salimans et al., 2024; Yan et al., 2024). Consistency models (CMs) (Song et al., 2023; Song and Dhariwal, 2023) introduce a self-consistency constraint and support standalone training as a new class of generative models. sCM (Lu and Song, 2024) extends this consistency constraint to continuous time to minimize discretization error. iMM (Zhou et al., 2025) incorporates moment matching. Shortcut (Frans et al., 2024) redefines the boundary condition to allow transport between arbitrary timesteps. MeanFlow (Geng et al., 2025) and AYF (Sabour et al., 2025) further extend Shortcut to continuous time. However, these methods still tend to produce slightly blurry results on large-scale text-to-image and video tasks (Luo et al., 2023a), so distributional matching methods, such as adversarial training (Lin et al., 2024; Ren et al., 2024; Lin and Yang, 2024; Lin et al., 2025a, b; Lu et al., 2025; Sauer et al., 2024b, a; Wang et al., 2024; Kohler et al., 2024; Kang et al., 2024; Xu et al., 2024; Chen et al., 2025) and score distillation (Yin et al., 2024b, a; Sauer et al., 2024b; Lu et al., 2025; Luo et al., 2023b; Zheng et al., 2025b), are often incorporated in practice.
@@
Early GAN research developed many techniques that succeeded on domain-specific datasets (Reed et al., 2016; Zhang et al., 2017; Karras et al., 2017, 2019, 2020b, 2021, 2020a). BigGAN (Brock et al., 2018) and StyleGAN-XL (Sauer et al., 2022) further scaled GANs to ImageNet (Russakovsky et al., 2015). However, GANs have fallen out of favor because of their training instability and limited scalability. Several works on large-scale text-to-image generation with GANs still employ convolutional architectures with complex designs (Kang et al., 2023; Zhu et al., 2025). GANs with transformer architectures have been challenging to scale (Jiang et al., 2021; Lee et al., 2021; Hudson and Zitnick, 2021). More recently, R3GAN (Huang et al., 2024) simplifies the adversarial formulation and achieves state-of-the-art performance on the ImageNet-64 benchmark using a convolutional architecture. GAT (Hyun et al., 2025) further extends this line of work to a latent transformer architecture. These works have revitalized interest in adversarial training. However, GAT still employs a non-standard transformer architecture and relies on a pre-trained feature network. Our work combines adversarial models with flow models and improves the training stability of adversarial methods.
@@ 3.1 Adversarial Training Preliminaries
Our method builds on generative adversarial networks (GANs), in which a generator $G:\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ aims to transport samples $z$ from a prior distribution $\mathcal{Z}$, e.g., a Gaussian, to samples $x$ from the data distribution $\mathcal{X}$, while a discriminator $D:\mathbb{R}^{n}\rightarrow\mathbb{R}$ aims to distinguish real samples from generated ones. The adversarial optimization involves a minimax game in which $D$ is trained to maximize this distinction, while $G$ is trained to minimize it:
$\mathcal{L}^{D}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x}\left[f(D(x),D(G(z)))\right],$ (1)
$\mathcal{L}^{G}_{\mathrm{adv}}$ $=\mathbb{E}_{z,x}\left[f(D(G(z)),D(x))\right].$ (2)
We adopt the relativistic objective (Jolicoeur-Martineau, 2018), where $f(a,b)=-\log(\mathrm{sigmoid}(a-b))$, because it yields a better loss landscape (Sun et al., 2020) and achieves the current state of the art (Huang et al., 2024; Hyun et al., 2025).
Additionally, gradient penalties $R_{1}$ and $R_{2}$ (Roth et al., 2017) are applied to $D$. They prevent $G$ from being pushed away from equilibrium (Mescheder et al., 2018) and impose a constraint on the Lipschitz constant of $D$ (Gulrajani et al., 2017). Directly computing these gradient penalties requires expensive double backpropagation and second-order differentiation, so we use a finite-difference approximation (Lin et al., 2025a):
$\mathcal{L}^{D}_{\mathrm{r1}}$ $=\mathbb{E}_{x}\left[\|\nabla_{x}D(x)\|^{2}_{2}\right]$ (3)
@@
$\mathcal{L}^{D}_{\mathrm{r2}}$ $=\mathbb{E}_{z}\left[\|\nabla_{G(z)}D(G(z))\|^{2}_{2}\right]$ (5)
@@
where $\epsilon$ is set to 0.01. We compute the penalties on only 25% of the samples in each batch and observe no performance degradation.
To prevent the discriminator logits from drifting unboundedly in the relativistic setting, we add a logit-centering penalty, following prior work (Karras et al., 2017):
$\small\mathcal{L}^{D}_{\mathrm{cp}}=\mathbb{E}_{z,x}\left[(D(x)+D(G(z)))^{2}\right].$ (7)
The final GAN objectives are:
@@
where $\lambda_{\mathrm{gp}}$ is a tuned hyperparameter that controls the scale of both gradient penalties, $R_{1}$ and $R_{2}$, and $\lambda_{\mathrm{cp}}$ is fixed at 0.01.
The expectations are estimated with Monte Carlo approximations over minibatches during training. The generator $G$ and discriminator $D$ are updated alternately. For conditional generation, the condition $c$ is provided to both networks as $G(z,c)$ and $D(x,c)$. we omit this notation when it is not needed.
@@ 3.2 Single-step Adversarial Flow Models
The GAN objective above only enforces that $G(z)$ matches the data distribution. However, there are many valid transport maps from $z$ to $x$, and the model is free to learn any one of them. Our proposed adversarial flow models instead learn a deterministic optimal transport map. This prevents generator drift and stabilizes training.
Formally, in optimal transport theory, Brenier’s theorem guarantees the existence of a unique optimal transport map when the source distribution is absolutely continuous, e.g., Gaussian, and the cost function is quadratic.
Accordingly, in adversarial flow models, we parameterize the transport map with a deterministic neural network $G(z)$. We further restrict the prior distribution to have the same dimensionality $n$ as the data distribution, i.e., $x,z\in\mathbb{R}^{n}$ and $G:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. This constraint is commonly required by flow-based models and does not reduce the generality of our method.
Our goal is to find an optimal transport map $G^{*}$ that is both a valid transport map and the map that minimizes the total transport cost under the quadratic cost function $\mathbf{c}(x,z)=\|x-z\|^{2}_{2}$. This corresponds to Wasserstein-2 optimal transport:
$\small G^{*}=\arg\min_{G}\int_{\mathcal{Z}}\mathbf{c}(G(z),z)\,dz.$ (10)
Since the adversarial objective encourages matching between the generated and target distributions, and achieves exact marginal matching at the global optimum under standard assumptions, e.g., infinite capacity and perfect optimization, the validity of the transport map is enforced. Under these assumptions, we find that an optimal transport regularization loss can be applied to $G$ to bias the solution toward the unique optimal transport map $G^{*}$. This loss minimizes the expectation of $\mathbf{c}(G(z),z)$ over $z$, which by definition minimizes the total transport cost:
$\boxed{\mathcal{L}^{G}_{\mathrm{ot}}=\mathbb{E}_{z}\left[\frac{1}{n}\|G(z)-z\|^{2}_{2}\right]}$ (11)
$\hskip 17.0pt=\frac{1}{n}\int_{\mathcal{Z}}\|G(z)-z\|^{2}_{2}\,dz.$ (12)
The objectives of adversarial flow models (AF) become:
@@
In practice, however, the validity of the transport map is not strictly enforced, because $\mathcal{L}^{G}_{\mathrm{ot}}$ competes with $\mathcal{L}^{G}_{\mathrm{adv}}$ and biases the solution toward $G(z)=z$. We therefore introduce a scaling factor $\lambda_{\mathrm{ot}}$. As illustrated in Figure 2, if this value is too small, optimization may get stuck in local minima; if it is too large, it pushes the model toward the identity map and hurts distribution matching. We therefore adopt a schedule that decreases $\lambda_{\mathrm{ot}}$ over the course of training.
@@
Figure 2: The effect of using different $\lambda_{\textrm{ot}}$ scales. $\lambda_{\textrm{ot}}=0$ is equivalent to GANs. $\lambda_{\textrm{ot}}$ being too small fails to escape local minima. $\lambda_{\textrm{ot}}$ being too large forces identity output.
Empirically, we observe that our method consistently learns the same deterministic mapping across different random initializations in the 1D mixture-of-Gaussians experiment (Figure 1), whereas standard GANs produce different transport mappings.
Unlike consistency-based methods, our model does not need to be trained on other timesteps of the probability flow and can instead be trained directly for one-step generation. This saves model capacity, reduces the number of training iterations, and avoids error propagation. Furthermore, in the one-step setting, this formulation removes the hyperparameters associated with timestep sampling and weighting. Our one-step model also avoids teacher forcing entirely.